OpenStudy (anonymous):

write expression. some number minus 18 is equal to the number times -20

OpenStudy (anonymous):

so far I have " y-18 =" I cant get the other part

OpenStudy (anonymous):

is it -20y?

Directrix (directrix):

Let n be the some number. Then, some number minus 18 is represented by n - 18 -------------------- Now, translate this: some number minus 18 is equal to the number times -20 some number minus 18 is equal to the number times -20 n - 18 = n*(-20) or n - 18 = -20*n ---> What about this? @teebabyy

OpenStudy (anonymous):

thank you!

Directrix (directrix):

Happy to help.